OpenStudy (anonymous):

s(t)= 1.85sin(36piet) t= seconds find velocity for t=0.025sec

OpenStudy (anonymous):

\[v(t)=1.85*36pisin(36\pi t)\] \[v(0.025)=1.85*36\pi \sin(36 \pi 0.025)\] pi = 3.14

OpenStudy (anonymous):

sorry!!!!

OpenStudy (anonymous):

\[v(0.025)=1.85*36 \pi \cos(36 \pi 0.025)\]

OpenStudy (anonymous):

??

OpenStudy (anonymous):

derivative of sin is cos

OpenStudy (anonymous):

what is your problem now??

OpenStudy (anonymous):

why is it(36*pie-cos)

OpenStudy (anonymous):

\[d(\sin(at))/dt=a \cos(at)\]

OpenStudy (anonymous):

so where do i go from there

OpenStudy (anonymous):

s(t) is a distance function, so the velocity is derivative of s(t) \[v(t)=ds(t)/dt\]

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Velocity

OpenStudy (anonymous):

how can i get my answer is cm/s

OpenStudy (anonymous):

Since your question doesnt involve any distances, you can choose your time interval to be 1 cm per 1 second (t).